OpenStudy (anonymous):

find the length of the diameter of a sphere with a volume of 3053.63ft^3

sam (.sam.):

Formula is \[V=\frac{4}{3}\pi r^3\] \[r=\sqrt[3]{\frac{3}{4\pi} V}\] 2 radius = diameter \[2r=d\] so \[\frac{d}{2}=\sqrt[3]{\frac{3}{4\pi} V}\] \[d=2\sqrt[3]{\frac{3}{4\pi} V}\]

OpenStudy (anonymous):

i got 48.5

OpenStudy (anonymous):

\[V={{4\over3}\pi r^3}\]solve for \(r\):\[{3053.63\times{3\over4\pi}}=r^3\]\[r^3=14263.53\]\[r\approx24.25\]\[d=2r\]\[d=2(24.25)\]\[d=...?\]

OpenStudy (anonymous):

yes that is correct :)\[d=48.5ft\]

OpenStudy (anonymous):

but the answer is not right it gotta be one of these 18ft, 10.9ft, 21.8ft, 9ft

OpenStudy (anonymous):

sorry i messed up the answer should be 38.61ft...either you wrote your question wrong or your answers are incorrect

sam (.sam.):

Careful, use your calculator \[\LARGE d=2(\sqrt[3]{\frac{3}{4\pi} V})\]

OpenStudy (anonymous):

i used a calculator, put in exactly that, and I got \(\approx38.61ft\)

OpenStudy (anonymous):

is it 21.8